OpenStudy (anonymous):

consider the function f(x) = {(sinx)/x, x cannot equal 0 {k, , x = 0 In order for f(x) to be continuous at x - 0, the value of k must be...

OpenStudy (anonymous):

We have that \[\lim_{x \rightarrow 0} (\sin x)/x =1\] (This is a common result; I'm sure you can find a proof of it somewhere on Google if you need to) Now in order for a function to be continuous at 0, we require that \[\lim_{x \rightarrow 0} f(x) =f(0)\] It's pretty clear that the only k for which this is true is k=1.

OpenStudy (anonymous):

sorry, just slightly confused on the end about the k

OpenStudy (blacksteel):

Zid's got it exactly right. In order for a function to be continuous, the value of the function at every point has to be equal to the value of the limit as x approaches that point. Thus, since \[\lim_{x \rightarrow 0} \sin(x)/x = 1\], we pick k to be 1 so that at x=0, f(x) = 1.

OpenStudy (anonymous):

ok, that makes sense.